# IWR-Bench Team

### v1

@@ 7 Contributors
Core Contributors ( Equal Contribution )
  • Yang Chen, Shanghai AI Lab, Zhejiang University
@@
  • Botian Shi, Shanghai AI Lab, Shanghai Innovation Institute
Project Leader
  • Yufan Shen, Shanghai AI Lab [Email: shenyfzju@gmail.com]
@@ Appendix A Multi-Dimensional Task Taxonomy
To ensure our benchmark reflects the breadth of real-world web applications, the third axis classifies tasks by their Application Domain. Drawing inspiration from established taxonomies in web-centric agent research (Gou et al., 2025), we group tasks into five high-level domains: Commerce & Services (e-commerce, booking, finance), Knowledge & Education (academic sites, news portals, documentation), Productivity & Tools (calculators, project management boards), Entertainment & Media (games, streaming platforms), and Lifestyle & Community (social forums, blogs). This classification guarantees that models are evaluated across a diverse spectrum of functionalities and visual paradigms. For instance, a task involving filtering products on an e-commerce site is tagged as [L2, V2, Commerce], while a task requiring the reconstruction of a simple browser game is tagged as [L4, V1, Entertainment]. This multi-dimensional labeling allows us to analyze whether a model’s performance correlates with certain types of interactions, visual styles, or application contexts.
@@ Appendix C Metric Parameter Validation
The VFS and Final Score metrics rely on the weighting coefficients $w$ and $\alpha$. To determine and validate these parameters, a human alignment study was conducted. A sample of 60 evaluation instances was constructed by selecting outputs from three randomly chosen models for each of 20 randomly selected tasks from IWR-Bench. Five PhD-level students assessed each instance on two dimensions: visual fidelity and overall quality. The optimal parameter values are determined by maximizing the Spearman’s $\rho$ correlation between the automated scores and the aggregated human judgments. For the VFS, peak correlation ($\rho=0.57$) is achieved with $w=0.5$, giving equal weight to LVS and HVS. For the Final Score, correlation with human overall judgment is maximized ($\rho=0.65$) with $\alpha=0.7$, empirically validating the decision to weigh functionality (IFS) more heavily than visual fidelity (VFS).

### v3

@@ 7 Contributors
Core Contributors***Equal contribution.
  • Yang Chen, Shanghai AI Lab, Zhejiang Universityzjucheny@gmail.com
@@
  • Yufan Shen, Shanghai AI Labshenyfzju@gmail.com
  • Botian Shi, Shanghai AI Lab, Shanghai Innovation Institute shibotian@pjlab.org.cn
Project Leader
@@ Appendix A Multi-Dimensional Task Taxonomy
To ensure our benchmark reflects the breadth of real-world web applications, the third axis classifies tasks by their Application Domain. Drawing inspiration from established taxonomies in web-centric agent research (gou2025mind2web2evaluatingagentic), we group tasks into five high-level domains: Commerce & Services (e-commerce, booking, finance), Knowledge & Education (academic sites, news portals, documentation), Productivity & Tools (calculators, project management boards), Entertainment & Media (games, streaming platforms), and Lifestyle & Community (social forums, blogs). This classification guarantees that models are evaluated across a diverse spectrum of functionalities and visual paradigms. For instance, a task involving filtering products on an e-commerce site is tagged as [L2, V2, Commerce], while a task requiring the reconstruction of a simple browser game is tagged as [L4, V1, Entertainment]. This multi-dimensional labeling allows us to analyze whether a model’s performance correlates with certain types of interactions, visual styles, or application contexts.
@@ Appendix C Metric Parameter Validation
The VFS and Final Score metrics rely on the weighting coefficients $w$ and $\alpha$. These parameters are determined and validated through a human alignment study. A sample of 60 evaluation instances is constructed by selecting outputs from three randomly chosen models for each of 20 randomly selected tasks from IWR-Bench. Each instance is assessed by five PhD-level students on two dimensions: visual fidelity and overall quality. To determine the optimal parameters, a grid search is performed over the discrete set $\{0.1,0.2,\dots,0.9\}$ for both $w$ and $\alpha$. The value that maximizes the Spearman’s $\rho$ correlation between the automated scores and the aggregated human judgments is selected. For the VFS metric, the peak correlation ($\rho=0.57$) is observed at $w=0.5$, indicating an equal weighting between LVS and HVS. For the Final Score, the maximum correlation with human overall judgment ($\rho=0.65$) is achieved with $\alpha=0.7$, empirically validating the decision to weigh functionality (IFS) more heavily than visual fidelity (VFS).